Fuzzy Monkey Technologies, Inc., purchased as a long-term investment $80 million of 8% bonds, dated January 1, on January 1, 2018. Management intends to have the investment available for sale when circumstances warrant. For bonds of similar risk and maturity the market yield was 10%. The price paid for the bonds was $66 million. Interest is received semiannually on June 30 and December 31. Due to changing market conditions, the fair value of the bonds at December 31, 2018, was $70 million. Required: 1. to 3. Prepare the relevant journal entries on the respective dates (record the interest at the effective rate). 4-a. At what amount will Fuzzy Monkey report its investment in the December 31, 2018, balance sheet? 4-b. Prepare the entry necessary to achieve this reporting objective. 5. How would Fuzzy Monkey's 2018 statement of cash flows be affected by this investment?
Answer #1
Debit Credit
Jun-30 Cash $       3.20
Discount on bond investment $       0.10
Interest revenue $       3.30
Dec-31 Cash $       3.20
Discount on bond investment $       0.11
Interest revenue $       3.31

4) 66 +0.10 +0.11 = 66.21

5)  Operating cash flow = 3.2 +3.2 =6.4  (inflow)

Investing cash flow = 66 ( Outflow)